OpenStudy (anonymous):

Find the solution of this equation. 7x-6=1

OpenStudy (anonymous):

I CAN

OpenStudy (anonymous):

7x-6=1 +6=1 you cancel out the 6 which now your left with 7x=7 you divide 7x/7 and get x and then 7/7 x=1 hope this help

OpenStudy (anonymous):

@arianadarsh good job :) you missed a "+6" on the right on the secont steps :P

OpenStudy (anonymous):

7x-6=1 7x=1+6 7x=7 x=7/7=1

OpenStudy (anonymous):

oh ya it hard to type it out it better to draw it but it comes out sloppy